OpenStudy (anonymous):

If cos theta = -5/13 and sin theta less than 0, what is tan theta?

OpenStudy (kittiwitti1):

Is this what you're asking?\[\text{If }\cos\theta=-\frac{5}{13}\text{ and }\sin\theta\lt0\text{, what is }\tan\theta?\]

OpenStudy (anonymous):

yessss

OpenStudy (kittiwitti1):

Okay. First question you ask yourself: which quadrant has negative sines?

OpenStudy (kittiwitti1):

(aka negative "y-values")

OpenStudy (anonymous):

okay um 1 and 3 i think

OpenStudy (kittiwitti1):

Q1 is wrong, but you got Q3. Here's a reference image: http://www.chilimath.com/algebra/intro/point/images/quadrants_1.gif

OpenStudy (anonymous):

i couldnt think without looking so 3 and 4

OpenStudy (kittiwitti1):

But good work nonetheless. =) So now you have narrowed down your options to Q3 and Q4.

OpenStudy (kittiwitti1):

(sorry I need to find some scratch paper)

OpenStudy (michele_laino):

hint angle theta has to belongs to the III quadrant: 

OpenStudy (michele_laino):

belong*

OpenStudy (kittiwitti1):

Okay yeah what she said. So now we have a triangle in QIII